# Picosecond Precision Heavy Ion Detector for Hypernuclei Lifetime Studies

### v3

@@ Abstract
In this paper, we present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetime. The detector employs the previously developed 10 picosecond resolution Radio Frequency (RF) Timer, which converts the temporal information of incident particles into spatial coordinates of secondary or photoelectrons on a position-sensitive detector by means of circular RF scanning in the 500–1000 MHz range. Here, we report the detector design to achieve efficient suppression of accidental background and effective separation of prompt reaction products and delayed events from hypernuclei decays, results of test studies carried out with RF synchronized laser as well as preliminary results obtained by using alpha particles. Dedicated Monte-Carlo simulations have been performed to estimate the detector’s performance under realistic experimental conditions at RF-driven electron, photon, or proton beams. The results confirm the feasibility of the proposed design and provide a basis for upcoming experimental measurements, based on the delayed fission detection.
@@ 1 Introduction
We present the design and preliminary performance evaluation of a new heavy ion detector for direct measurements of heavy hypernuclei lifetime at RF-driven electron, photon or proton beams. The system integrates the Radio Frequency Timer (RFT) as its core timing device and has been optimized to suppress accidental background and enable efficient separation of prompt and delayed events.
2 Heavy Ion Detector Concept and Design
@@
The Radio Frequency Heavy Ion Detector (RF HID) presented here builds on this RFT principle for applications in nuclear physics, specifically for direct lifetime measurements of heavy (A200) hypernuclei. Figure 1 illustrates its operational setup. Primary electron, photon, or proton bunches, driven by the RF system of the accelerator, hit the target (1) and produce secondary electrons (SE). Reaction products from beam particles, such as prompt or delayed fission fragments (FF) can also generate SEs, at a level of tens of SEs per fragment. The SEs emitted from both sides of the target are accelerated by an electrical potential (2.5 kV) applied between a cathode (target) and an accelerating electrode (2). They are deflected by 90∘ using a permanent magnet (3), collimated (4), focused by an electrostatic lens (5) and circularly scanned by the deflector (6) driven by the RF synthesizer (7) synchronized with the accelerator’s RF system (11). Coordinates of the deflected electrons are measured by the PSD (10).
@@
By selecting an appropriate position-sensitive detector (PSD), the system can be made selectively insensitive to prompt SEs, which arrive at a fixed time with respect to the beam bunch, and hence strike a specific position on the PSD. A 100 ps length shielding, which corresponds to 18-degree sector (figure 2) for 500 MHz scanning case, could be placed in front of the PSD. This shield will absorb the majority of promptly produced SEs and prevent the MCPs from rapid degradation. The lifetime distribution will then be accurately reconstructed from the detected low rate delayed events that fall outside the shielded phase region. The reference zero-time point can be determined under low beam intensity conditions and subsequently shifted, using a phase shifter, to the center of the shielded sector.
@@ 3 Laboratory Tests
A dedicated setup, employing one arm of the detector, was built in the laboratory for preliminary experimental studies of the detector’s performance. To test the detector’s response to secondary electrons generated by fission fragments, an alpha-particle source (239Pu) was used in place of the target. The emitted 5 MeV alpha particles pass through gold foil and produce SEs, which are then detected by the system. A 1 mm diameter collimator was employed (see section 5 for details on the collimation effects on resolution and efficiency). The resulting image of the scanned SEs is shown in figure 3. The elliptical image of detected electrons is clearly distinguishable above MCP background noise.
Additional tests were performed using RF-synchronized photons to characterize the system’s time resolution under controlled conditions. These measurements were carried out as part of previous photoemission lifetime studies of 2D materials [19]. The general layout of the setup is shown in figure 4. It includes an RFT, a NKT Origami [3] femtosecond laser synchronized with a 500-1000 MHz RF generator, feeding the RF deflector after amplification. The laser delivers 166 fs pulses at a wavelength of 515.44 nm, which is then halved to 257.7 nm using a frequency-doubling crystal. The photon pulses hit the target material and generate photoelectrons, which are subsequently accelerated, deflected and timed by the RFT.
Depending on the material under study, the intrinsic photoemission time may vary. Figure 5 (adapted from [19]) (top) shows the hit positions of the circularly scanned photoelectrons (left) and the corresponding phase distribution, converted into time units (right) for a 0.6 µm layer of gold. In this case, the time spread is about 12 ps, which matches the detector’s intrinsic resolution; therefore, these photoelectrons can be considered instant or “prompt”.
Notably, measurements with a monolayer graphene photoemitter revealed a time distribution closely resembling the expected lifetime spectrum of heavy hypernuclei decays. A significant number of prompt photoelectrons were detected, while smaller part of the events exhibited delays of up to 2 nanoseconds. These extended photoemission times in graphene can be theoretically attributed to its high optical phonon energy and limited phase space for acoustic phonon scattering, which result in hot carriers lifetimes ranging from hundreds of picoseconds to a few nanoseconds [12, 13]. Experimentally, the nanosecond-scale photoemission times in graphene at room temperature were first observed using a highly sensitive time-of-flight angle-resolved photoemission spectroscopy (TOF ARPES) technique operated at very low laser fluence (10 J/cm2) [13].
@@
For the estimation of prompt and delayed event rates, we considered the following configuration: 1.5 GeV electrons delivered in 500 MHz bunches with a duration of 2 ps irradiate a thin bismuth foil of thickness 4 mg/cm2, corresponding to T=0.00064 radiation lengths. At these beam energies and for such a thin target, electron–nucleus interactions are dominated by virtual-photon processes. Electron straggling due to bremsstrahlung can therefore be neglected. The yield Y of a photoproduction process with cross section (k), induced by a single electron through a virtual photon of energy k, is
@@
and, considering only the first reaction, the delayed-fission cross section for bismuth is roughly 12.5 b (see [29] and references therein).
Assuming prompt and delayed photo-fission cross sections of 2.5 mb and 12.5 b, respectively, in the 1.0–1.5 GeV energy range, a flux of 1012 electrons traversing a 4 mg/cm2 Bi target is expected to produce, on average, one delayed-fission event and roughly 200 prompt-fission events. Each fission event emits two oppositely directed fragments which, upon exiting the target, generate more than 100 secondary electrons per fragment (360 according to [16]). In contrast, a 1.5 GeV beam electron produces only 0.03 secondary electrons in the forward direction and 0.015 in the backward direction [1].
For a beam intensity of 1010 e/s, each 2 ns period delivers a 2 ps bunch containing 20 electrons, generating approximately 0.6 forward and 0.3 backward secondary electrons. A beam of this intensity yields about 200 prompt-fission events and at least one delayed-fission event within 100 s, with each fission event producing more than 100 secondary electrons on each side of the target. The time distribution of these secondary electrons is recorded with the HID system. Additionally, by using the PSD design described in section 2 (see figure 2), the detector can be made effectively insensitive to prompt events, allowing only the low-rate delayed candidate hypernuclear-production events to be registered and processed by the data acquisition system.
@@ 5 Expected lifetime measurement performance: Monte-Carlo simulations
Simulations based on SIMION 8 software package [6] were conducted to evaluate the detector’s performance under varying geometric configurations and applied static voltages, resulting in different combinations of detection efficiency, focusing distance, and time resolution. The schematic of the simulated setup is shown in figure 6. Secondary electrons (SEs) are generated with uniform spatial and angular distributions at emission area on target (1). The size of the emission area corresponds to the cross-section of the high-energy beam hitting a target. The initial energies of secondary electrons emitted by fission fragments have been experimentally measured to lie in the range of 0.3 to 5 eV [7, 16].
In the present simulations, the SE energy distribution is taken as exponential, with a mean value of 2 eV. The electrons are pre-focused. and accelerated to 2500 eV in a Wehnelt-type three-electrode system (2), deflected by 90 degrees in a permanent magnet (3), and then pass through the focusing system composed of a collimator (4) and two Einzel lenses (5) and (6), arranged sequentially. The transit time spread (TTS) and electron beam spot size at the detection plane (7) are calculated. The TTS represents the detector’s physical time resolution, while the spot size contributes to the technical time resolution, given by:
where T is the period of the RF cycle, d is the convolution of the SE spot size at the detection plane and the position resolution of the PSD, and r is the radius of the scanned circle. In the simulations we consider T = 2 ns (corresponding to a scan frequency 500 MHz), d is taken as the SE spot size at the detection plane (the spot size is in the hundreds of micron range, which is significantly larger than the position resolution of the PSD) and r = 10 mm (typical scan radius for 25 mm diameter MCPs). Figure 7 shows the calculated components of time resolution: SE transit time spread (), technical time resolution (), and total time resolution (), given by:
@@
Using different collimator aperture we can choose an optimal efficiency and detector resolutions. This calculations show that even in the worst scenario we can achieve 30ps resolution. Simulations also demonstrate that the secondary electron transport system can be positioned at distances up to 1 m or more from the target, if needed to mitigate radiation damage to the PSD and associated electronics, without significant degradation of time resolution or detection efficiency.
@@
For the simulation estimates presented below, we assume and prompt-to-delayed ratios 103 and 105. The measured prompt events times will follow a Gaussian distribution around the time zero point with a standard deviation determined by the detector’s time resolution, while the delayed events will follow an exponential distribution superimposed with Gaussian noise (again with standard deviation determined by the detector’s time resolution).
@@ 6 Summary
A new time-resolved heavy ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system can be used for the heavy ion precise timing and is particularly suitable for heavy hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot carrier lifetime in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.

### v4

# Picosecond Precision Heavy-Ion Detector for Hypernuclei Lifetime Studies

@@ Abstract
In this paper, we present the design and preliminary performance evaluation of a new heavy-ion detector for direct measurements of heavy hypernuclei lifetime. The detector employs the previously developed 10 picosecond resolution Radio Frequency (RF) Timer, which converts the temporal information of incident particles into spatial coordinates of secondary or photoelectrons on a position-sensitive detector by means of circular RF scanning in the 500–1000 MHz range. Here, we report the detector design to achieve efficient suppression of accidental background and effective separation of prompt reaction products and delayed events from hypernuclei decays, results of test studies carried out with RF synchronized laser as well as preliminary results obtained by using alpha particles. Dedicated Monte-Carlo simulations have been performed to estimate the detector’s performance under realistic experimental conditions at RF-driven electron, photon, or proton beams. The results confirm the feasibility of the proposed design and provide a basis for upcoming experimental measurements, based on the delayed fission detection.
@@ 1 Introduction
We present the design and preliminary performance evaluation of a new heavy-ion detector for direct measurements of heavy hypernuclei lifetime at RF-driven electron, photon or proton beams. The system integrates the Radio Frequency Timer (RFT) as its core timing device and has been optimized to suppress accidental background and enable efficient separation of prompt and delayed events.
2 Heavy-Ion Detector Concept and Design
@@
The Radio Frequency Heavy-Ion Detector (RF HID) presented here builds on this RFT principle for applications in nuclear physics, specifically for direct lifetime measurements of heavy (A200) hypernuclei. Figure 1 illustrates its operational setup. Primary electron, photon, or proton bunches, driven by the RF system of the accelerator, hit the target (1) and produce secondary electrons (SE). Reaction products from beam particles, such as prompt or delayed fission fragments (FF) can also generate SEs, at a level of tens of SEs per fragment. The SEs emitted from both sides of the target are accelerated by an electrical potential (2.5 kV) applied between a cathode (target) and an accelerating electrode (2). They are deflected by 90∘ using a permanent magnet (3), collimated (4), focused by an electrostatic lens (5) and circularly scanned by the deflector (6) driven by the RF synthesizer (7) synchronized with the accelerator’s RF system (11). Coordinates of the deflected electrons are measured by the PSD (10).
@@
By selecting an appropriate position-sensitive detector (PSD), the system can be made selectively insensitive to prompt SEs, which arrive at a fixed time with respect to the beam bunch, and hence strike a specific position on the PSD. A 100 ps length shielding, which corresponds to 18-degree sector (figure 2) for the 500 MHz scanning case, could be placed in front of the PSD. This shield will absorb the majority of promptly produced SEs and prevent rapid degradation of the MCPs. The lifetime distribution will then be accurately reconstructed from the detected low rate delayed events that fall outside the shielded phase region. The reference zero-time point can be determined under low beam intensity conditions and subsequently shifted, using a phase shifter, to the center of the shielded sector.
@@ 3 Laboratory Tests
A dedicated setup, employing one arm of the detector, was built in the laboratory for preliminary experimental studies of the detector’s performance. To test the detector’s response to secondary electrons generated by fission fragments, an alpha particle source (239Pu) was used in place of the target. The emitted 5 MeV alpha particles pass through a gold foil and produce SEs, which are then detected by the system. A 1 mm diameter collimator was employed (see Section 5 for details on the collimation effects on resolution and efficiency). The resulting image of the scanned SEs is shown in Fig. 3. The elliptical image of detected electrons is clearly distinguishable from MCP background noise.
Additional tests were performed using RF-synchronized photons to characterize the system’s time resolution under controlled conditions. These measurements were carried out as part of previous photoemission lifetime studies of 2D materials [19]. The general layout of the setup is shown in Fig. 4. It includes an RFT, a NKT Origami [3] femtosecond laser synchronized with a 500-1000 MHz RF generator, feeding the RF deflector after amplification. The laser delivers 166 fs pulses at a wavelength of 515.44 nm, which is then halved to 257.7 nm using a frequency-doubling crystal. The photon pulses hit the target material and generate photoelectrons, which are subsequently accelerated, deflected and timed by the RFT.
Depending on the material under study, the intrinsic photoemission time may vary. Figure 5 (adapted from [19]) (top) shows the hit positions of the circularly scanned photoelectrons (left) and the corresponding phase distribution, converted into time units (right) for a 0.6 µm layer of gold. In this case, the time spread is about 12 ps, which matches the detector’s intrinsic resolution; therefore, these photoelectrons can be considered instantaneous or “prompt”.
Notably, measurements with a monolayer graphene photoemitter revealed a time distribution closely resembling the expected lifetime spectrum of heavy hypernuclei decays. A significant number of prompt photoelectrons were detected, while a smaller part of the events exhibited delays of up to 2 nanoseconds. These extended photoemission times in graphene can be theoretically attributed to its high optical phonon energy and limited phase space for acoustic phonon scattering, which result in hot-carrier lifetimes ranging from hundreds of picoseconds to a few nanoseconds [12, 13]. Experimentally, the nanosecond-scale photoemission times in graphene at room temperature were first observed using a highly sensitive time-of-flight angle-resolved photoemission spectroscopy (TOF ARPES) technique operated at very low laser fluence (10 J/cm2) [13].
@@
For the estimation of prompt and delayed event rates, we considered the following configuration: 1.5 GeV electrons delivered in 500 MHz bunches with a duration of 2 ps irradiate a thin bismuth foil of thickness 4 mg/cm2, corresponding to T=0.00064 radiation lengths. At these beam energies and for such a thin target, electron–nucleus interactions are dominated by virtual-photon processes. Electron straggling due to bremsstrahlung can therefore be neglected. The yield Y of a photon-induced reaction channel with cross section (k), where (k) denotes the cross section of the specific process under consideration, can be written in the general form:
@@
In a photoproduction experiment on a bismuth target performed at the Thomas Jefferson National Accelerator Facility, correlations between the produced protons, pions, and kaons and the fission of the recoil nucleus were investigated. The cross section for the formation of heavy hypernuclei associated with positive-kaon production (corresponding to only the first reaction in Eq. 4.2) was estimated to be approximately 12.5 b (see [29] for a detailed discussion).
Assuming prompt and delayed photo-fission cross sections of 2.5 mb and 12.5 b, respectively, in the 1.0–1.5 GeV energy range, 1012 electrons traversing a 4 mg/cm2 Bi target are expected to produce, on average, one delayed-fission event and roughly 200 prompt-fission events. Each fission event emits two oppositely directed fragments which, upon exiting the target, generate more than 100 secondary electrons per fragment (360 according to [16]). In contrast, a 1.5 GeV beam electron produces only 0.03 secondary electrons in the forward direction and 0.015 in the backward direction [1].
For a beam intensity of 1010 e/s, each 2 ns period delivers a 2 ps bunch containing 20 electrons, generating approximately 0.6 forward and 0.3 backward secondary electrons. A beam of this intensity yields about 200 prompt-fission events and at least one delayed-fission event within 100 s, with each fission event producing more than 100 secondary electrons on each side of the target. The time distribution of these secondary electrons is recorded with the HID system.
Additionally, by using the PSD design described in section 2 (see figure 2), the detector can be made effectively insensitive to prompt events, allowing only the low-rate delayed candidate hypernuclear-production events to be registered and processed by the data acquisition system.
@@ 5 Expected lifetime measurement performance: Monte-Carlo simulations
Simulations based on the SIMION 8 software package [6] were conducted to evaluate the detector’s performance under varying geometric configurations and applied static voltages, resulting in different combinations of detection efficiency, focusing distance, and time resolution. The schematic of the simulated setup is shown in Fig 6. Secondary electrons (SEs) are generated with uniform spatial and angular distributions at emission area on target (1). The size of the emission area corresponds to the cross-section of the high-energy beam hitting a target. The initial energies of secondary electrons emitted by fission fragments have been experimentally measured to lie in the range of 0.3 to 5 eV [7, 16].
In the present simulations, the SE energy distribution is taken as exponential, with a mean value of 2 eV. The electrons are pre-focused and accelerated to 2500 eV in a Wehnelt-type three-electrode system (2), deflected by 90 degrees in a permanent magnet (3), and then pass through the focusing system composed of a collimator (4) and two Einzel lenses (5) and (6), arranged sequentially. The transit time spread (TTS) and electron beam spot size at the detection plane (7) are calculated. The TTS represents the detector’s physical time resolution, while the spot size contributes to the technical time resolution, given by:
where T is the period of the RF cycle, d is the convolution of the SE spot size at the detection plane and the position resolution of the PSD, and r is the radius of the scanned circle. In the simulations we consider T = 2 ns (corresponding to a scanning frequency of 500 MHz), d is taken as the SE spot size at the detection plane (the spot size is in the hundreds of micron range, which is significantly larger than the position resolution of the PSD) and r = 10 mm (typical scan radius for 25 mm diameter MCPs). Figure 7 shows the calculated components of time resolution: SE transit time spread (), technical time resolution (), and total time resolution (), given by:
@@
By varying the collimator aperture we can choose an optimal efficiency and detector resolutions. These calculations show that even in the worst scenario we can achieve 30ps resolution. Simulations also demonstrate that the secondary-electron transport system can be positioned at distances up to 1 m or more from the target, if needed to mitigate radiation damage to the PSD and associated electronics, without significant degradation of time resolution or detection efficiency.
@@
For the simulation estimates presented below, we assume and prompt-to-delayed ratios 103 and 105. The measured prompt events times follow a Gaussian distribution around the time zero point with a standard deviation determined by the detector’s time resolution, while the delayed events follow an exponential distribution convoluted with Gaussian noise (again with standard deviation determined by the detector’s time resolution).
@@ 6 Summary
A new time-resolved heavy-ion detector, based on circular radio-frequency scanning of secondary electrons, has been designed, built and tested in the laboratory using a synchronized femtosecond laser. The measured time resolution of the detector is approximately 12 ps. This system can be used for the heavy-ion precise timing and is particularly suitable for heavy hypernuclear lifetime studies, enabling the separation of high-rate prompt reaction products from delayed events of interest. Laboratory studies of hot-carrier lifetime in graphene, which exhibit a time structure similar to that expected for hypernuclear decays, demonstrate the detector’s capability for such measurements. Additionally, the detection of secondary electrons generated by alpha particles has been successfully verified. Detailed Monte Carlo simulations confirm that the expected statistical uncertainty for direct hypernuclear lifetime measurements with this detector is about 10 ps under realistic conditions. These results indicate that the proposed detector concept is feasible and provides a solid basis for future experiments with electron, photon, or proton beams.